OpenStudy (anonymous):

Which of the following objects is in equilibrium? A. an object that moves at constant acceleration B. an object that moves at constant velocity C. an object that was thrown straight upward at the moment it had zero velocity D. an object moving in a circle at constant speed **Not sure :( Thank you!!

OpenStudy (danjs):

hello

OpenStudy (danjs):

Equilibrium means zero net external forces are acting on the object

OpenStudy (anonymous):

yes :)

OpenStudy (danjs):

An object that is accelerating must have a force by newtons famous second law, F=m*a

OpenStudy (danjs):

When an object is thrown straight up in a gravity field, the entire time it is experiencing a force due to gravity

OpenStudy (danjs):

Even when it is at rest at the top, there still is gravitational force pulling it down

OpenStudy (anonymous):

so based off of that, it would be C? thrown straight upward the moment it has 0 velocity?

OpenStudy (danjs):

an object moving in a circle has to have some force moving it in that circle, directed towards the center of the circle

OpenStudy (anonymous):

because nothing is acting on it?

OpenStudy (danjs):

Otherwise it would move in a straight line

OpenStudy (danjs):

No, there is always gravitational force accelerating the ball downwards

OpenStudy (anonymous):

ohh okay, so that means C and D are both out? it's between A and B? because C has gravitational force and the spinning in a circle (D) has some kind of force as well?

OpenStudy (danjs):

Now imagine you were moving at a constant velocity, F=m*a a=0 (constant velocity) so F = 0

OpenStudy (anonymous):

okay:)

OpenStudy (danjs):

did you talk about reference frames and inertial observers in class?

OpenStudy (anonymous):

ermm i believe so, but i need to refresh on that

OpenStudy (danjs):

okay, post up more questions if ya need !

OpenStudy (danjs):

Read Richard Feynman's books, you will love physics after that, probably

OpenStudy (anonymous):

okay, so wait, for this, the solution is actually D you said? where it spins?

OpenStudy (danjs):

No, an object on a curved trajectory has an acceleration towards the center of the circle

OpenStudy (danjs):

an acceleration can loosely be defined as , speed up , slow down, or change in direction

OpenStudy (anonymous):

and okay, i will see if my library has them:) ohh so it's the one moving in a straight line? constant velocity? so B?

OpenStudy (danjs):

yes, correct B

OpenStudy (anonymous):

ah okay, awesome! :) thank you:)

OpenStudy (danjs):

yw